OpenStudy (anonymous):

Find the derivative of G(x) if G(x) = the definite integral of(2lnydy)

OpenStudy (anonymous):

d..integral = ylny - y ; derivative g(x) = y'lny+ylny'-y'

OpenStudy (anonymous):

what? how do i actually write that

OpenStudy (anonymous):

u = lny; dv = 1 -> u'=1/y, v=y ....

OpenStudy (anonymous):

what? lol

OpenStudy (anonymous):

waiting ? i speak a little english .. no explain

OpenStudy (anonymous):

nevermind

OpenStudy (anonymous):

intergral partial product ?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

idk ?

OpenStudy (anonymous):

are u solve?

OpenStudy (anonymous):

no